What is the pH of a mixture of 5 ml of 0.1 mol/litre sodium acetate and 4 ml of 0.1 mol/litre acetic acid? (pKa of acetic acid at 25oC = 4.76)


1
Expert's answer
2021-08-11T11:22:01-0400

pKa=4.76

Ka=1.74*10-5

"Ka = \frac {[H^+] [CH_3COO^-]} {CH_3COOH]}"

"[H^+] = \frac {1.74 × 10^ {-5} × 0.0005}{0.0004} = 2.175 × 10^ {-5}"

"pH=-log_{10}(2.175\u00d710^{-5})=4.66"